TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Dan Dubrick
date: 2003-05-24 14:51:00
subject: 5\13 Pt 1 ISS On-Orbit Status 13-05-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

14 May 2003

ISS On-Orbit Status 13 May 2003

Part 1 of 2

All ISS systems continue to function nominally, except as noted
previously or below.

After their wakeup (2:00am EDT), the crew was commended on their good
job with yesterday's emergency response training session, which bodes
well for the fire response exercise scheduled later this week.

"Telephone and Telegraph Day" on board:  The crew performed a major
3-hr. routine maintenance job on the SM's STTS telephone/telegraph
subsystem and its audio paths and comm stations.  As part of the
activity, Malenchenko and Lu also tested VHF receivers, conducted an
inventory audit of headsets and cables with push-to-talk units, and
finally switched the STTS to its backup string for report to MCC-M.
[The "Voskhod-M" STTS enables telephone communications between the
SM, FGB, DC-1 and USOS (US segment), and also with users on the
ground over VHF channels selected by an operator at an SM comm panel,
via STTS antennas on the outside of the SM.  There are six comm
panels in the SM with pushbuttons for accessing any of three audio
channels plus an intercom channel.  Other modes of the STTS include
telegraphy (teletype), EVA voice, emergency alarms, Packet, and TORU
docking support.] 

CDR Yuri Malenchenko conducted a test activation and checkout of the
AD blood pressure monitor which he replaced yesterday in the Gamma-1M
medical complex. [For the checkout, he donned the PKO biomedical
harness with the blood pressure cuff (fastened over the left upper
arm) and its pressure hose, then connected the harness to the PMO
medical examination panel for blood pressure and rheoplethysmograph
readings.] 

The CDR completed the daily routine maintenance on the SOZh life
support system, including ASU toilet facility, KRP food containers,
EDV water containers and KTO solid waste containers.

Yuri and Ed performed their daily physical exercise program, on TVIS
(treadmill with vibration isolation and stabilization), RED
(resistive exercise device), CEVIS (cycle ergometer with vibration
isolation) and VELO cycle ergometer with load trainer.

Ed Lu was thanked by the ground for the great job performed yesterday
on the TVIS inspection.  A few engineering questions were uplinked to
clarify some observations.

Malenchenko and FE/SO Ed Lu today again had an hour reserved on their
schedule for ISS familiarization and adjustment to their new
surroundings and activities.

The crew was provided with the Expedition 6 post-landing debriefing
video. 

Today's CEO (crew earth observations) targets, severely restricted by
limited the use of the science window in the Lab due to the current
XPOP attitude, were Anatahan volcano, N Marianas (this island
volcano, captured by the crew last Sunday, may still be erupting.
Looking left 1-2 degrees off track.  Anatahan lies ~200 miles north
of Guam), Rome, Italy (left of track, inland about 20 miles from the
coast), Lisbon, Portugal (nadir pass. Sunny weather after the passage
of a very slow storm), Detroit, Michigan (nadir pass.  First clear
pass after many days of poor weather), Havana, Cuba (looking slightly
left), Guadalajara, Mexico (looking slightly left), Angolan Biomass
Burning (dry-season burning already under way in the most flammable
part of the globe, Angola's thick savanna forests.  Looking left for
views of individual fires as well the smoke pall), Industrialized SE
Africa smog (looking left for possible smog buildup over the
plateau.  Then nadir views of mountainous Lesotho, Africa's
Switzerland, at nadir), Tuamotu Islands, South Pacific (nadir views
of the west end of this island chain. Looking for fringing coral
reefs around numerous atolls.  This was the first opportunity to
shoot these islands in many weeks). http://eol.jsc.nasa.gov and
http://earthobservatory.nasa.gov 

U.S. and Russian Segment Status (as of 2:06pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control
(TCS):
Elektron O2 generator is powered On (16 Amp mode).  Vozdukh CO2
scrubber is On (manual mode 3).  U.S. CDRA CO2 scrubber is Off.
TCCS (trace contaminant control subsystem) is operating.  MCA (major
constituents analyzer) is operating.  BMP Harmful Impurities unit:
absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air
conditioner SKV-1 is On; SKV-2 is Off. SM Working Compartment:
Pressure (mmHg) -- 742; temperature (deg C) -- 28.5; ppO2 (mmHg) --
data invalid; ppCO2 (mmHg) -- data invalid. SM Transfer Compartment:
Pressure (mmHg) -- 751; temperature (deg C) -- 20.9.
FGB Cabin:  Pressure (mmHg) -- 752; temperature (deg C) -- 23.0.
Node:  Pressure (mmHg) -- 745.16; temperature (deg C) -- 23.5 (shell);
ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
U.S. Lab:  Pressure (mmHg) -- 747.14; temperature (deg C) -- 22.3;
ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
Joint Airlock (Equip. Lock):  Pressure (mmHg) -- 747.24; temperature
(deg C) -- 23.1; shell heater temp (deg C) -- 23.1, ppO2 (mmHg) --
172.5; ppCO2 (mmHg) -- 4.2.
PMA-1:  Shell heater temp (deg C) -- 24.6
PMA-2:  Shell heater temp (deg C) -- 12.3

(n/a = data not available)

Propulsion System (PS):

Total propellant load available: 3588 kg (7910 lb) as of 5/8
[SM(774) + FGB(2562) + Progress(252)].  (Capability: SM -- 860 kg;
FGB -- 6120 kg). 

Electrical Power Systems (EPS):

Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B
and 4B both in solar-tracking Auto mode ("sun slicer" position).
SM batteries:  Battery #4 is disconnected; battery #8 suspected
failed; battery #1 in "Cycle" mode; all other batteries (5) are in
"Partial Charge" mode.
FGB batteries:  Battery #1 is in "Cycle" mode; battery #4 is in ROM
(capacity restoration) mode, in parallel with #1; all other
batteries (4) are in "Partial Charge" mode.
Plasma Contactor Unit PCU-1 and PCU-2 both in Standby mode.

Command & Data Handling Systems:

C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
GNC-1 MDM is prime; GNC-2 is Backup.
INT-1 is operating; INT-2 is Off.
EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-2 MDM is On (primary); PL-1 MDM is Off
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
Being upgraded.
SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.
Being upgraded.

Attitude Control Systems:

3 CMGs on-line (CMG-1 failed).
State vector source -- U.S. SIGI-1 (GPS)
Attitude source -- U.S. SIGI-2 (GPS)
Angular rate source -- RGA-2

 - Continued -

@Message posted automagically by IMTHINGS POST 1.30
--- 
* Origin: SpaceBase(tm) Pt 1 -14.4- Van BC Canada 604-473-9358 (1:153/719.1)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.